# Revisiting XL-MIMO Channel Estimation: When Dual-Wideband Effects Meet Near Field

### v3

@@ III-B Angular-Delay Representation of SnS XL-MIMO Channel
Building on approaches used for traditional massive MIMO channels [9], the wideband XL-MIMO channels in (13) can be approximated as
@@
$i_{l,s}=\psi_{l}-2\varphi_{l}{n_{l,s}},\quad i_{l,e}=\psi_{l}-2\varphi_{l}{n_{l,e}}.$ (17)
@@
$\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\int_{i_{l,e}}^{i_{l,s}}c(\epsilon_{l})\mathbf{b}_{\mathrm{far}}(\epsilon_{l})\mathrm{d}\epsilon_{l},$ (18)
where $c(\epsilon_{l})$ denotes the decomposition coefficient, and $[\mathbf{b}_{\mathrm{far}}(\epsilon_{l})]_{n}=\mathrm{e}^{\mathrm{j}2\pi\epsilon_{l}n}$ represents the far-field array response corresponding to spatial frequency $\epsilon_{l}\in[i_{l,e},i_{l,s}]$. Furthermore, the continuous decomposition in (18) can be discretized with a spatial frequency resolution of $2/N$ as
@@
where the number of sampling points is given by $P=\lceil({n_{l,s}}-{n_{l,e}})\varphi_{l}N\rceil$, and the discretized spatial frequency is defined as $\epsilon_{l,p}=\min(i_{l,s},i_{l,e})+{2(p-1)}/{N}$ for $p=\{1,2,\dots,P\}$. The coefficient $c_{l,p}$ denotes the corresponding weight associated with the $p$-th far-field array response component.
@@
Similar to the analysis in (17), the start and end spatial frequencies of the SnS array response vector $\boldsymbol{\theta}(f_{k},\psi_{l},\varphi_{l})$ can be approximated as
$i_{l,s}(k)$ $=(a_{k}-b_{k}n_{s}){\lambda_{\mathrm{c}}}/{d},$ (21)
$i_{l,e}(k)$ $={(a_{k}-b_{k}n_{e})}\lambda_{\mathrm{c}}/{d}.$ (22)
@@
Lemma 1. By utilizing equations (20)-(22), it can be shown that the two-dimensional inverse discrete Fourier transform (2D-IDFT) of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ results in a sparse matrix, where the nonzero entries are confined within a square region. Within this region, the dominant components exhibit a clear column-wise clustered sparsity structure. Specifically, as $N,K\rightarrow\infty$, the support of the nonzero region converges to:
$[\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{l},\\ 0,&(n,k)\notin\mathcal{A}_{l},\end{cases}$ (24)
@@
$\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,e}\leq k\leq J_{l,s}\right\},$ (25)
@@
$J_{l,e}$ $=\max_{n}(t_{l,n}+{2}/{K}+1/2)K,$
$J_{l,s}$ $=\min_{n}(t_{l,n}-{2}/{K}+1/2)K,$
indicating the boundaries.
In addition, within the region $\mathcal{A}_{l}$, each column corresponds to the angular response associated with a specific propagation delay. The width of the significant spatial frequency components is primarily determined by the frequency-width effect. Assuming the starting spatial frequency of the $k$-th column is denoted by $i_{l,k}\in[i_{l,e}^{\min},i_{l,s}^{\max}]$, the corresponding angular spread can be approximated as
$B_{l,k}=\frac{i_{l,k}f_{s}}{f_{c}}N.$ (27)
As a result, the significant components of $\boldsymbol{\Xi}_{l}$ within $\mathcal{A}_{l}$ are confined to narrow bands along each column, thereby exhibiting a pronounced column-wise clustered sparsity pattern.
To summarize, by leveraging (24)-(27), the sparse structure of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ can be accurately characterized. In particular, equation (24) delineates the overall boundaries of the nonzero region, while (27) provides a more precise description of the angular spread in each column, further refining the support of the dominant components.
@@
$\mathcal{B}_{l,p}$ =$=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right.$ (28)
@@
  • Angular Spread: First, the spherical wavefront effect and the SnS characteristics determine the angular spread range at each subcarrier, characterized by $i_{l,s}(k)$ and $i_{l,e}(k)$. Then, the frequency-wideband effect governs the overall boundaries of the angular spread, denoted by $I_{l,e}$ and $I_{l,s}$. Moreover, the frequency-wideband effect also enables a more accurate description of the angular spread within each column, represented by $B_{l,k}$.
  • Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, leading to delay spread. The SnS properties define the range of variation for $n$, while the spherical wavefront imposes a quadratic trend on the delay spread as a function of $n$.
In summary, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively govern the angular and delay spreads, thereby reshaping the sparsity structure of SnS dual-wideband channels in the angular-delay domain. Specifically,
@@
To enhance intuitive understanding of the SnS XL-MIMO channel, we provide an example with the following parameters: $N=K=1024$, $f_{s}=3$GHz, $f_{c}=30$GHz, $r_{l}=10$m, $\vartheta_{l}=\pi/5$, and $\boldsymbol{\phi}_{l}\in[64,959]$. Fig. 2 illustrates the sparsity pattern of the SnS XL-MIMO channel in the angular-delay domain. Given the specified parameters, we have $i_{l,e}(0)=0.64$, $i_{l,s}(0)\approx 0.8$, $j_{l,s}(n_{s})=513$, and $j_{l,e}(n_{s})\approx 517$. Similarly, $i_{l,s}(K-1)\approx 0.88$, $i_{l,e}(K-1)\approx 0.72$, $j_{l,s}(n_{e})=545$, and $j_{l,e}(n_{e})\approx 549$. As depicted in Fig. 2(a) and Fig. 2(b), these values provide accurate boundaries for the ranges of angular and delay spread. Additionally, the figures show that both the angular and delay spectra shift to the right due to $\cos\vartheta_{l}>0$ and the monotonicity of $t_{n}$ with respect to $n$.
Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,e}=513$, $J_{l,s}=549$, and $B_{k,s}\approx 41$ with $k=J_{l,e}$ Fig. 2(c) illustrates the non-zero region of $\mathbf{F}_{\mathrm{A}}^{{}_{\mathrm{H}}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}^{*}_{\mathrm{D}}$, which is accurately constrained to $\mathcal{A}_{l}$, validating the effectiveness of Lemma 1. In particular, for the $J_{l,e}$-th column, the width of angular spread also aligns the result of (27), which further validates the influence of frequency-wideband effect for local angular spread.
@@
Compared to conventional massive MIMO-OFDM channels, the proposed SnS dual-wideband XL-MIMO channel exhibits distinct behavior. Specifically, in addition to beam squint, the spherical wavefront effect and SnS properties contribute to angular spread, resulting in broader angular diffusion. Moreover, the quadratic phase variations introduced by the spherical wavefront reshape the trend of delay variations across antennas. Notably, Lemma 1 is applicable to conventional MIMO-OFDM channels. When the spherical wavefront simplifies to a plane wavefront and SnS properties are disregarded, both $\varphi_{l}$ and $\kappa_{l,n}$ can be omitted. Thus, we have $i_{l,e}(k)=i_{l,s}(k)$ and $t_{l,n}=nf_{s}\psi_{l}$. Consequently, the significant components are uniformly distributed within the non-zero region, as illustrated in [9].
@@ IV Problem Formulation and Sparse Prior Modeling
In this section, leveraging the sparsity outlined in Theorem 1, we first formulate the SnS dual-wideband XL-MIMO channel estimation problem as a sparse signal recovery task. Subsequently, to account for the global block sparsity and local common-delay sparsity, we introduce a novel column-wise sparse prior model tailored to this structure.
@@ IV-A Problem Formulation
Motivated by the sparsity of SnS dual-wideband XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a MMV-based sparse recovery problem. Utilizing the angular-delay representation in (14), the received signal model in (3) can be further written as
@@
While this work primarily focuses on the ULA configuration, the signal models in (29) and (30) are also applicable to the uniform planar array (UPA) configuration. Specifically, consider a BS equipped with a half-wavelength UPA comprising $N=N_{v}\times N_{h}$ antennas, where $N_{v}$ and $N_{h}$ denote the number of elements along the vertical and horizontal dimensions, respectively. In this case, the angular transformation matrix $\mathbf{F}_{\mathrm{A}}$ can be redefined as the Kronecker product $\mathbf{F}_{\mathrm{A}}=\mathbf{F}_{\mathrm{A},v}\otimes\mathbf{F}_{\mathrm{A},h}$, where $\mathbf{F}_{\mathrm{A},v}\in\mathbb{C}^{N_{v}\times N_{v}}$ and $\mathbf{F}_{\mathrm{A},h}\in\mathbb{C}^{N_{h}\times N_{h}}$ denote the vertical and horizontal DFT matrices, respectively. It is worth noting that the sparsity structure in the angular-delay domain may differ between ULA and UPA configurations, particularly with regard to local common-delay sparsity, due to the inherent differences in array geometry and two-dimensional angular resolution. Although a comprehensive analysis of the UPA case is beyond the scope of this paper, we acknowledge its importance and leave it as a promising direction for future work.
In this paper, our primary objective is to develop an effective estimation algorithm for accurately reconstructing $\mathbf{X}$ or $\mathbf{H}$ according to $\mathbf{Y}$ and $\boldsymbol{\Psi}$. In the existing literature, several sparse signal recovery algorithms have been employed to address the problem in (30), including SOMP [30, 36] and optimization-based methods [37, 38]. However, these methods fail to effectively exploit the inherent sparsity structure of the angular-delay channel and often require additional parameters, such as the rank of $\mathbf{X}$ or the number of multipaths, which may not always be available. Considering the block sparsity of $\mathbf{X}$, Bayesian inference techniques such as approximate message passing (AMP) and SBL have shown superior recovery performance in channel estimation tasks [39, 40, 41]. However, the effectiveness of these Bayesian methods relies heavily on the accuracy of the prior model.
@@ IV-B Column-Wise Hierarchical Sparse Prior Model
According to Lemma 1 and Theorem 1, the angular-delay domain channel exhibits prominent block sparsity and a column-wise clustered structure. However, existing sparse prior models fail to capture these characteristics fully [40, 39, 42, 41]. For example, the two-layer hierarchical prior models in [40, 39] assign independent priors to the variance precision parameters of the channel coefficients, which limits their ability to model the correlation across coefficients. Although [42, 41] leverage MRF to promote global block sparsity, they still overlook local structural dependencies, especially those induced by column-wise clustering. As a result, these models are not well suited for the SnS wideband XL-MIMO channel estimation problem considered in this work.
To this end, we propose a novel column-wise hierarchical sparse prior model, which incorporates both precision parameter sharing mechanism and MRF structure to promote structured sparsity. Specifically, the column-wise hierarchical prior is defined as
@@
1) Variance-Driven Sparsity: The first layer adopts a complex Gaussian prior with zero mean and variance $[\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$, i.e.,
@@
where $\gamma_{n,k}$ denotes the variance precision parameter, with $\gamma_{n,k}^{-1}$ representing the variance of $x_{n,k}$. This precision-based modeling naturally induces sparsity: as the variance $\gamma_{n,k}^{-1}$ approaches zero, the corresponding $x_{n,k}$ is effectively pushed toward zero, thereby promoting a sparse solution.
2) Precision Parameters Sharing Mechanisms: Due to the finite angular and delay resolutions, the angular-delay domain channel does not exhibit strict sparsity, as the coefficients near the boundaries of the sparse patterns often retain non-negligible energy. Enforcing hard zeros on these components may lead to degraded estimation performance. To mitigate this issue, in the second layer, we adopt a dual-precision parameter strategy, wherein each variance $\gamma_{n,k}$ is associated with two distinct precision parameters corresponding to its active and inactive states. In addition, to capture the local common-delay clustered sparsity, we further introduce the precision parameters sharing mechanism. Specifically, $\gamma_{n,k}$ is assigned a conditionally Bernoulli-Gaussian distribution given by
@@
As described in (33) and (34), this common-delay sparsity mechanism ensures that, for a fixed $k$, all $x_{n,k}$ sharing the same variance state are governed by a common variance precision parameter. This coupling enforces a consistent sparsity structure among components associated with the same delay.
@@
where $\mathcal{D}{n,k}$ denotes the set of neighboring nodes of $s_{n,k}$; $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$ and $v(s_{n,k})=\exp(-\eta s_{n,k})$ represent the pairwise potential and self-potential functions, respectively, with $\varpi$ and $\eta$ being the model parameters associated with $p(\mathbf{S})$. By leveraging the formulation in (35), the prior encourages block patterns in the support structure and suppresses isolated coefficients that deviate from their neighboring states.

### v4

@@ III-B Angular-Delay Representation of SnS XL-MIMO Channel
To enable efficient channel estimation, leveraging channel sparsity is critical. Regarding the spatial domain, the Euclidean and polar domains are theoretically preferable transformation choices where each propagation path in these domains may correspond to a single line spectrum, yielding a highly sparse representation[21]. However, these domains rely on joint two-dimensional sampling, which leads to transformation matrices with drastically high dimensionality, and consequently, heightened computational complexity.
As an alternative, we propose leveraging angular-delay sparsity for estimation [9]. Thus, the spatial-frequency channels in (13) can be approximated as
@@
$\mathcal{I}_{l,s}=\psi_{l}-2\varphi_{l}{n_{l,s}},\quad\mathcal{I}_{l,e}=\psi_{l}-2\varphi_{l}{n_{l,e}}.$ (17)
@@
$\boldsymbol{\kappa}_{l}\odot\mathbf{b}(\psi_{l},\varphi_{l})=\int_{\mathcal{I}_{l,e}}^{\mathcal{I}_{l,s}}c(\epsilon_{l})\mathbf{b}_{\mathrm{far}}(\epsilon_{l})\mathrm{d}\epsilon_{l},$ (18)
where $c(\epsilon_{l})$ denotes the decomposition coefficient, and $\mathbf{b}_{\mathrm{far}}(\epsilon_{l})]_{n}=\mathrm{e}^{\mathrm{j}2\pi\epsilon_{l}n}$ represents the far-field array response corresponding to spatial frequency $\epsilon_{l}\in[\mathcal{I}_{l,e},\mathcal{I}_{l,s}]$. Furthermore, the continuous decomposition in (18) can be discretized with a spatial frequency resolution of $2/N$ as
@@
where the number of sampling points is given by $P=\lceil({n_{l,s}}-{n_{l,e}})\varphi_{l}N\rceil$, and the discretized spatial frequency is defined as $\epsilon_{l,p}=\min(\mathcal{I}_{l,s},\mathcal{I}_{l,e})+{2(p-1)}/{N}$ for $p=\{1,2,\dots,P\}$. The coefficient $c_{l,p}$ denotes the corresponding weight associated with the $p$-th far-field array response component.
@@
Similar to the analysis in (17), the start and end spatial frequencies of the frequency-dependent phase-shift vector $\boldsymbol{\theta}(f_{k},\psi_{l},\varphi_{l})$ can be approximated as
$i_{l,s}(k)$ $=(a_{l,k}-b_{l,k}n_{l,s}){\lambda_{\mathrm{c}}}/{d},$ (21)
$i_{l,e}(k)$ $={(a_{l,k}-b_{l,k}n_{l,e})}\lambda_{\mathrm{c}}/{d}.$ (22)
@@
Lemma 1. By utilizing (20)-(22), it can be shown that the two-dimensional inverse discrete Fourier transform (2D-IDFT) of $\boldsymbol{\Theta}(\psi_{l},\varphi_{l})$ results in a sparse matrix, where the nonzero entries are confined within a square region. Within this region, the dominant components exhibit a clear column-wise clustered sparsity structure. Specifically, as $N,K\rightarrow\infty$, the support of the nonzero region converges to:
$\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}]_{n,k}=\begin{cases}\text{nonzeros},&(n,k)\in\mathcal{A}_{l},\\ 0,&(n,k)\notin\mathcal{A}_{l},\end{cases}$ (24)
@@
$\mathcal{A}_{l}\triangleq\left\{(n,k)\in\mathbb{Z}^{2}\mid I_{l,e}\leq n\leq I_{l,s},J_{l,s}\leq k\leq J_{l,e}\right\},$ (25)
@@
$J_{l,s}$ $=\min_{n}j_{l,s}(n),\,J_{l,e}=\max_{n}j_{l,e}(n),$
indicating the boundaries with $j_{l,s}(n)=(t_{n}-{2}/{K}+1/2)K$ and $j_{l,e}(n)=(t_{n}+{2}/{K}+1/2)K$.
@@
Within the region $\mathcal{A}_{l}$, each column corresponds to the angular response associated with a specific propagation delay. The width of the significant spatial frequency components is primarily determined by the frequency-width effect. Assuming the starting spatial frequency of the $k$-th column is denoted by $i_{l,k}\in[\min_{k}i_{l,e}(k),\max_{k}i_{l,s}(k)]$, the corresponding angular spread can be approximated as
$B_{l,k}=\frac{i_{l,k}f_{s}}{f_{c}}N.$ (27)
As a result, the significant components of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{\mathrm{*}}$ within $\mathcal{A}_{l}$ are confined to narrow bands along each column, thereby exhibiting a pronounced column-wise clustered sparsity pattern. To summarize, by leveraging (24)-(27), the sparse structure of $\mathbf{F}_{\mathrm{A}}^{\mathrm{H}}\boldsymbol{\Theta}(\varphi_{l},\psi_{l})\mathbf{F}_{\mathrm{D}}^{*}$ can be accurately characterized. In particular, (24) determines the overall boundaries of the nonzero region, while (27) provides a more precise description of the angular spread in each column, further refining the support of the dominant components.
@@
$\mathcal{B}_{l,p}$ $=\left\{(n,k)\in\mathbb{Z}^{2}\mid n=\mathrm{mod}(n_{l}+N\epsilon_{l,p},N),\right$. (28)
@@
  • Angular Spread: First, the spherical wavefront effect and the SnS characteristics determine the angular spread range at each subcarrier, characterized by $i_{l,s}(k)$ and $i_{l,e}(k)$. Then, the frequency-wideband effect governs the overall boundaries of the angular spread, denoted by $I_{l,e}$ and $I_{l,s}$, and enables a more accurate description of the angular spread within each column, represented by $B_{l,k}$.
  • Delay Spread: The spatial-wideband effect causes variations in $j_{l,e}(n)$ and $j_{l,s}(n)$ across antennas $n$, which in turn leads to delay spread. The SnS properties define the range of variation for the antenna index $n$, while the spherical wavefront imposes a quadratic trend on the delay spread as a function of $n$.
To summarize, the spherical wavefront effect, SnS properties, and dual-wideband effects collectively govern the angular and delay spreads, thereby reshaping the sparsity structure of XL-MIMO channels in the angular-delay domain. Specifically,
@@
To enhance intuitive understanding of the XL-MIMO channel, we provide an example with the following parameters: $N=K=1024$, $f_{s}=3$GHz, $f_{c}=30$GHz, $r_{l}=10$m, $\vartheta_{l}=\pi/5$, and $\boldsymbol{\phi}_{l}\in[64,959]$. Fig. 2 illustrates the sparsity pattern of the SnS XL-MIMO channel in the angular-delay domain. Given the specified parameters, we have $i_{l,e}(0)=0.64$, $i_{l,s}(0)\approx 0.8$, $j_{l,s}(n_{s})=513$, and $j_{l,e}(n_{s})\approx 517$. Similarly, $i_{l,e}(K-1)\approx 0.72$, $i_{l,s}(K-1)\approx 0.88$, $j_{l,s}(n_{e})=545$, and $j_{l,e}(n_{e})\approx 549$. As depicted in Fig. 2(a) and Fig. 2(b), these values provide accurate boundaries for the ranges of angular and delay spread.
Furthermore, we obtain the following values: $I_{l,e}=841$, $I_{l,s}=962$, $J_{l,s}=513$, $J_{l,e}=549$, and $B_{k,s}\approx 41$ with $k=J_{l,e}$. Fig. 2(c) illustrates the non-zero region of $\mathbf{F}_{\mathrm{A}}^{{}_{\mathrm{H}}}\boldsymbol{\Theta}(\psi_{l},\varphi_{l})\mathbf{F}^{*}_{\mathrm{D}}$, which is accurately constrained to $\mathcal{A}_{l}$, validating the effectiveness of Lemma 1. In particular, for the $J_{l,s}$-th column, the width of angular spread also aligns the result of (27), which further validates the influence of frequency-wideband effect for local angular spread.
@@
Compared to massive MIMO-OFDM channels, the wideband XL-MIMO channel exhibits distinct behavior. Specifically, in addition to beam squint, the spherical wavefront effect and SnS properties contribute to angular spread, resulting in broader angular spread. Moreover, the quadratic phase variations introduced by the spherical wavefront reshape the trend of delay variations across antennas. Notably, Lemma 1 and Theorem 1 are applicable to massive MIMO-OFDM channels. When the spherical wavefront simplifies to a plane wavefront and SnS properties are disregarded, both $\varphi_{l}$ and $\kappa_{l,n}$ can be omitted. Thus, we have $i_{l,e}(k)=i_{l,s}(k)$ and $t_{l,n}=nf_{s}\psi_{l}$. Consequently, the significant components are uniformly distributed in the non-zero region [9].
@@ IV Problem Formulation and Sparse Prior Modeling
In this section, leveraging the sparsity outlined in Theorem 1, we first formulate the XL-MIMO channel estimation problem as a sparse recovery task. Subsequently, to capture the global block sparsity and local common-delay sparsity, we introduce a column-wise sparse prior model.
@@ IV-A Problem Formulation
Motivated by the sparsity of XL-MIMO channels in the angular-delay domain, the channel estimation problem can be formulated as a MMV-based sparse recovery problem. Utilizing the angular-delay representation in (14), the received signal model in (3) can be further written as
@@
While this work primarily focuses on the ULA configuration, the signal model in (30) is also applicable to the uniform planar array (UPA) configuration. Specifically, consider a half-wavelength UPA comprising $N=N_{v}\times N_{h}$ antennas, where $N_{v}$ and $N_{h}$ denote the number of elements along the vertical and horizontal dimensions, respectively. In this case, the angular transformation matrix $\mathbf{F}_{\mathrm{A}}$ can be redefined as $\mathbf{F}_{\mathrm{A}}=\mathbf{F}_{\mathrm{A},v}\otimes\mathbf{F}_{\mathrm{A},h}$, where $\mathbf{F}_{\mathrm{A},v}\in\mathbb{C}^{N_{v}\times N_{v}}$ and $\mathbf{F}_{\mathrm{A},h}\in\mathbb{C}^{N_{h}\times N_{h}}$ denote the vertical and horizontal DFT matrices, respectively.
This paper aims to develop an effective estimation algorithm for reconstructing $\mathbf{H}$ according to $\mathbf{Y}$ and $\boldsymbol{\Psi}$. So far, several sparse signal recovery algorithms have been employed to address the problem in (30), including SOMP [29, 36] and optimization-based methods [37, 38]. However, these methods fail to effectively exploit the inherent sparsity structure of the angular-delay channel and often require additional knowledge, such as the rank of $\mathbf{X}$ or the number of multipaths, which may not always be available. Considering the block sparsity of $\mathbf{X}$, Bayesian inference techniques such as approximate message passing (AMP) and SBL have shown superior recovery performance in channel estimation tasks [39, 40, 41]. However, the effectiveness of these Bayesian methods relies heavily on the accuracy of the prior model.
@@ IV-B Column-Wise Hierarchical Sparse Prior Model
According to Lemma 1 and Theorem 1, the angular-delay domain channel exhibits prominent block sparsity and a column-wise clustered sparsity structure. However, existing sparse prior models fail to capture these characteristics fully [40, 39, 42, 41]. For example, the two-layer hierarchical prior models in [40, 39] assign independent priors to the precision parameters of the channel coefficients, which limits their ability to model the correlation across coefficients. Although [42, 41] leverage MRF to promote global block sparsity, they still overlook local structural dependencies, especially those induced by column-wise clustering. As a result, these models are not well suited for the XL-MIMO channel estimation problem considered in this work.
To this end, we propose a novel column-wise hierarchical sparse prior model, which incorporates both precision sharing mechanism and MRF structure to promote structured sparsity. Specifically, the column-wise hierarchical prior is defined as
@@
1) Variance-Driven Sparsity: In the first layer, we adopt a complex Gaussian prior with mean zero and variance $\boldsymbol{\Gamma}]_{n,k}=\gamma_{n,k}^{-1}$ for each coefficient, i.e.,
@@
where $\gamma_{n,k}^{-1}$ denotes the variance corresponding to coefficient $x_{n,k}$. This variance-driven modeling naturally induces sparsity. Specifically, as the variance approaches zero, the corresponding $x_{n,k}$ is effectively pushed toward zero, thereby promoting a sparse solution.
2) Precision Sharing Mechanism: Considering the sparsity of angular-delay channel, we adopt a dual-precision strategy, where each $\gamma_{n,k}$ is associated with two distinct precision parameters corresponding to the active and inactive states of channel coefficients. Building on this, to capture the local common-delay clustered sparsity, we further introduce the precision sharing mechanism, where $\gamma_{n,k}$ is assigned a conditionally Bernoulli-Gamma distribution as
@@
As described in (33) and (34), this common-delay sparsity mechanism ensures that, for a fixed $k$, all $x_{n,k}$ sharing the same state are governed by a common precision parameter. This coupling enforces a consistent clustered sparsity structure among components associated with the same delay, which aligns with the local common-delay sparsity.
@@
where $\mathcal{D}_{n,k}$ denotes the set of neighboring nodes of $s_{n,k}$; $u(s_{n,k},s_{n^{\prime},k^{\prime}})=\exp(\varpi s_{n,k}s_{n^{\prime},k^{\prime}})$ and $v(s_{n,k})=\exp(-\eta s_{n,k})$ represent the pairwise potential and self-potential functions, respectively, with $\varpi$ and $\eta$ being the model parameters associated with $p(\mathbf{S})$. By leveraging the formulation in (35), the prior encourages block patterns in the support structure and suppresses isolated coefficients that deviate from their neighboring states.